# SpatialRGPT: Grounded Spatial Reasoning in Vision-Language Models

### v2

@@ 5 Conclusion
Appendix A More Results on Standard Visual-langauge Benchmarks
In this study, we present results on MM-VET, as shown in Table 3. Our method significantly improves spatial reasoning tasks compared to baselines that are not trained with OSD. These results highlight that even without local region prompts, our framework can still enhance spatial reasoning.
rec ocr know gen spatial math total
RegionVILA 37.1 24.9 21.2 19.6 26.4 3.5 33.3
SpatialRGPT 34.2 24.7 17.5 17.4 28.9 3.8 31.5
SpatialRGPT-Depth 36.1 27.3 20.5 17.7 32.1 7.7 33.9
Table 3: Spatial RGPT Results on MMVet.
Appendix B Statistics and Samples of SpatialRGPT-Bench
Figure 7: Samples in SpatialRGPT-Bench.
Appendix C Ablation Study on Different Input Modalities
@@
SpatialRGPT-Mask 99.17 99.04 80.19 91.96 87.50 91.81 91.78
SpatialRGPT-Box 99.17 98.09 83.01 91.96 82.69 92.72 91.47
@@
SpatialRGPT-Mask 41.2 / 0.33 65.6 / 0.25 51.9 / 0.27 49.6 / 0.31 57.9 / 0.61 95.3 / 15.4°
SpatialRGPT-Box 39.2 / 0.35 63.1 / 0.25 56.6 / 0.27 48.8 / 0.36 60.1 / 1.06 94.3 / 10.2°
@@
Appendix D Implementation Details for Data Pipeline
@@
D.2 Metric Depth Estimation
@@
D.3 Point Cloud Processing
@@
D.4 Open Spatial Dataset QA Templates
We provide a few samples for each category of QA in the templates that we use to generate QAs mentioned in Section 3.1. Due to space limitations, we are unable to list the full set here. We plan to release our code to the public upon publication.
@@ D.5 LLM Prompts for Complex QA
Appendix E Implementation Details for SpatialRGPT Architecture
E.1 Visual Backbone.
We adopt a pre-trained OpenAI CLIP-L model [65] as the visual backbone. We use $\times×336$ image resolutions to include more visual details for the model, which can help with vision language tasks that require fine-grained details [49] and are beneficial for region-level representations [33].
E.2 Region-feature Extractor.
@@ Appendix E Implementation Details for SpatialRGPT Architecture
E.3 Multi-modal Connector
@@
Appendix F Implementation Details for Training SpatialRGPT
F.1 Instruction Tuning Data
@@ F.2 Hyperparameters
Appendix G Experimental Setting and Details
G.1 Experiments Compute Resources
@@
Appendix H Benchmark Evaluation Details
@@
Appendix I Limitations and Future Work
For the most accurate object detection, oriented bounding boxes (OBB) are preferred over axis-aligned bounding boxes (AABB). As illustrated in Figure 10, the dimensions obtained from AABBs can differ from those obtained with OBBs. There are two methods to compute an OBB. A simple method involves calculating the OBB using Principal Component Analysis (PCA) of the object’s convex hull, which provides an approximate minimal bounding box. However, this approximation often lacks accuracy due to the incomplete and noisy nature of point clouds captured from a single view. Furthermore, this method still cannot handle extreme cases when objects are partially elevated. The most precise method involves determining the OBB based on the object’s pose, which is currently challenging due to limitations in obtaining accurate object poses. Future improvements could include integrating available pose estimation approaches. However, currently, there are no open-vocabulary solutions for object pose estimation, so this remains an area for future research.
@@ Appendix I Limitations and Future Work
Appendix J Broader Impacts

### v3

@@ Conclusion.
Limitations.
One limitation of our work is the use of Axis-Aligned Bounding Boxes (AABBs), which can result in inaccuracies in label representation. A more accurate alternative is oriented bounding boxes (OBBs), but implementing them requires precise object pose estimation, which remains challenging due to the lack of open-world solutions. The most accurate approach would be human labeling [74], although this requires significant effort. We leave these for future work.
Acknowledgement.
This work was supported, in part, by the Qualcomm Innovation Fellowship.
Appendix A Ablation Study on Augmented SpatialRGPT-Bench
We conduct additional experiments by augmenting and rephrasing both questions and answers in SpatialRGPT-Bench using GPT-4. The results are shown in Table 5. The results show that SpatialRGPT consistently outperforms the baseline models, even when the questions and answers are different from the training data.
Below/
Above
Left/
Right
Big/
Small
Tall/
Short
Wide/
Thin
Behind/
Front
Qualitative
Average
GPT-4V-Turbo 66.7 47.6 66.0 64.2 71.1 47.2 60.5
SpatialRGPT-7B 95.8 99.0 77.4 92.9 82.7 90.9 90.0
Direct
Distance
Horizontal
Distance
Vertical
Distance
Width
Height
Direction
GPT-4V-Turbo 30.4 / 0.87 26.2 / 2.66 33.9 / 0.51 48.8 / 0.35 69.1 / 1.35 40.1 / 70.0°
SpatialRGPT-7B 43.2 / 0.32 63.9 / 0.27 52.8 / 0.26 51.1 / 0.31 54.1 / 1.02 95.3 / 15.3°
Table 5: Augmented SpatialRGPT-Bench results. Numbers represent success rates ($\uparrow$) and absolute relative error ($\downarrow$).
Appendix B Ablation Study on Metric-Scale Width and Height Data
We conduct an ablation study to see if adding width and height data affects other types of questions. As shown in Table 6, adding this data slightly improved the accuracy for questions about size (like big/small, tall/short, wide/thin) but slightly worsened the accuracy for questions about the distance between objects (horizontal and vertical). This suggests that information about object size helps with size-related questions but might make distance measurements less clear.
Below / Above
Left / Right
Big / Small
Tall / Short
Wide / Thin
Behind / Front
Avg.
- width & height 99.1 99.0 75.8 90.8 82.8 92.1 90.5
+ width & height 99.1 +0   99.0 +0   80.1 +4.3   91.9 +1.1   87.5 +4.7   91.8 -0.3   90.5 +1.2
Direct Distance
Horizontal Distance
Vertical Distance
Width
Height
Direction
- width & height 41.2 69.3 54.8 22.8 21.2 95.1
+ width & height 41.2 +0   65.6 -3.7   51.9 -2.9   49.6 +26.8   57.9 +36.7   95.3 +0.2
Table 6: Ablation study on the impact of width and height data on the performance of other categories. Numbers represent success rates ($\uparrow$).
Appendix C Ablation Study on Bounding Box Types
We conduct an ablation study to examine the effect of using axis-aligned bounding boxes (AABB) versus PCA-based oriented bounding boxes (OBB). For this study, we use human-labeled OBBs from the Omni3D test set as the ground truth. We then compare the mean-square error of the width and height measurements for AABBs and PCA-based OBBs labeled by our 3D scene graph pipeline. The results are shown in Table 7. PCA-based OBB often lacks accuracy due to the incomplete and noisy nature of point clouds captured from a single view.
BBox Type Width (↓) Height (↓)
Oriented BBox 17.09 4.83
Axis-aligned BBox 8.27 2.35
Table 7: Ablation study on axis-aligned vs. oriented bounding boxes. Numbers indicate MSE comparing to Omni3D ground truth.
Appendix D Ablation Study on Different Input Modalities
@@
SpatialRGPT-7B-Mask 99.17 99.04 80.19 91.96 87.50 91.81 91.78
SpatialRGPT-7B-Box 99.17 98.09 83.01 91.96 82.69 92.72 91.47
@@
SpatialRGPT-7B-Mask 41.2 / 0.33 65.6 / 0.25 51.9 / 0.27 49.6 / 0.31 57.9 / 0.61 95.3 / 15.4°
SpatialRGPT-7B-Box 39.2 / 0.35 63.1 / 0.25 56.6 / 0.27 48.8 / 0.36 60.1 / 1.06 94.3 / 10.2°
@@
Appendix E Statistics and Samples of SpatialRGPT-Bench
Figure 8: Samples in SpatialRGPT-Bench.
Appendix F Implementation Details for Data Pipeline
@@
F.2 Metric Depth Estimation
@@
F.3 Point Cloud Processing
@@
F.4 Open Spatial Dataset QA Templates
We provide samples for each category of QA in the templates that we use to generate QAs mentioned in Section 3.1.
@@ F.5 LLM Prompts for Complex QA
Appendix G Implementation Details for SpatialRGPT Architecture
G.1 Visual Backbone.
We adopt a pre-trained OpenAI CLIP-L model [64] as the visual backbone. We use 336$\times$336 image resolutions to include more visual details for the model, which can help with vision language tasks that require fine-grained details [50] and are beneficial for region-level representations [34].
G.2 Region-feature Extractor.
@@ Appendix G Implementation Details for SpatialRGPT Architecture
G.3 Multi-modal Connector
@@
Appendix H Implementation Details for Training SpatialRGPT
H.1 Instruction Tuning Data
@@ H.2 Hyperparameters
Appendix I Experimental Setting and Details
I.1 Experiments Compute Resources
@@
Appendix J Benchmark Evaluation Details
@@
Appendix K More Discussion on Limitations
For the most accurate object detection, oriented bounding boxes (OBB) are preferred over axis-aligned bounding boxes (AABB). As illustrated in Figure 11, the dimensions obtained from AABBs can differ from those obtained with OBBs. There are two methods to compute an OBB. A simple method involves calculating the OBB using Principal Component Analysis (PCA) of the object’s convex hull, which provides an approximate minimal bounding box. However, this approximation often lacks accuracy due to the incomplete and noisy nature of point clouds captured from a single view. Furthermore, this method still cannot handle extreme cases when objects are partially elevated (see Appdx C). The most precise method involves determining the OBB based on the object’s pose, which is currently challenging due to limitations in obtaining accurate object poses. Future improvements could include integrating available pose estimation approaches. However, currently, there are no open-vocabulary solutions for object pose estimation, so this remains an area for future research. Another direction, explored in subsequent work (e.g., Q-Spatial Bench [74]), addresses this limitation by leveraging human labeling.
@@
Appendix L Broader Impacts